# FALCON: Fine-grained Activation Manipulation by Contrastive Orthogonal Unalignment for Large Language Model

### v1

@@ 4.2 Contrastive Orthogonal Unalignment
To achieve selective knowledge unlearning in LLMs, we devise Contrastive Orthogonal Unalignment through contrastive mechanisms and gradient projection to balance knowledge unlearning and retention.
@@ 4.2.1 Contrastive Representation Unlearning
To enhance unlearning, we construct Principal Offset Vectors (POV), which steer model activations away from undesired knowledge by reducing their alignment with dominant components identified via SVD in the representation space. The goal of POVs is guiding activations into less entangled subspaces, effectively separating the forget and retain sets.
@@
$\mathcal{H}^{+}=\frac{f\left(r\cdot\left(I-w\sum_{i=1}^{K}v_{i}v_{i}^{\top}\right),\epsilon\right)}{\left|f\left(r\cdot\left(I-w\sum_{i=1}^{K}v_{i}v_{i}^{\top}\right),\epsilon\right)\right|}$ (8)
Here, $r\in\mathbb{R}^{D}$ is a vector initialized with random values, $w$ controls the influence of principal directions, and $I\in\mathbb{R}^{D\times D}$ is the identity matrix. The term $\epsilon$ introduces optional perturbations. The function $f(\cdot)$ represents a generalized transformation operator like projection and nonlinear mapping, designed to enhance disentanglement and make representations more resistant to recovery.
This formulation ensures $\mathcal{H}^{+}$ is directed away from dominant principal subspaces, combining deterministic guidance and randomized transformations to improve robustness. Unlike generic random vectors, POVs target multiple dominant features, enhancing resistance to adversarial prompts and improving unlearning effectiveness.
For each input sample, we define three types of representations: the anchor representation $\mathcal{H}_{a}$ from the updated model for the forget set, the positive representation $\mathcal{H}^{+}$, given by the POV defined in Equation 8, and the negative representations $\mathcal{H}^{-}$ from the frozen model. To ensure consistent scaling, all representations are normalized, and their similarity scores are measured using cosine similarity:
$S^{+}$ $=\sum_{d=1}^{D}\mathcal{H}_{a}[d]\cdot\mathcal{H}^{+}[d]$
$S^{-}$ $=\sum_{z=1}^{\mathcal{Z}}\sum_{d=1}^{D}\mathcal{H}_{a}[d]\cdot\mathcal{H}_{z}^{-}[d]$ (9)
@@
$\mathcal{L}_{\mathcal{F}}=-\frac{1}{|B|}\sum_{b=1}^{|B|}\log\frac{\exp(S_{b}^{+}/\tau)}{\exp(S_{b}^{+}/\tau)+\sum_{b=1}^{\mathcal{N}}\exp(S_{b}^{-}/\tau)}$ (10)
where $\tau$ is a temperature scaling parameter. This loss encourages the updated model’s representations to align with the Principal Offset Vectors while diverging from the frozen model’s representations of undesired knowledge. By leveraging both directional guidance through Principal Offset Vectors and contrastive learning, our approach achieves more precise and efficient representation unalignment in activation space.
In addition to unlearning undesired representations, it is essential to preserve critical knowledge required for downstream tasks. To achieve this, we define a retain loss that measures the alignment between the updated model’s activations ($\mathcal{H}^{u}$) and the frozen model’s activations ($\mathcal{H}^{f}$) for the retain set. This alignment is evaluated using a retention alignment loss, which can be viewed as a self-supervised variant of contrastive loss. By maximizing the alignment consistency between the updated and frozen model activations, this loss ensures that critical knowledge is retained effectively during the unlearning process. The retain loss $\mathcal{L}_{\mathcal{R}}$ is defined as:
@@
4.2.2 Orthogonalizing Gradient Conflict
After computing the forget loss $\mathcal{L}_{\mathcal{F}}$ and retain loss $\mathcal{L}_{\mathcal{R}}$, we address the misalignment between the optimization directions of unlearning and retaining by employing a gradient projection mechanism. This mechanism orthogonalizes conflicting gradients onto a subspace, minimizing interference and promoting balanced optimization.
Given the gradients of the forget and retain losses, denoted as $\nabla\mathcal{L}_{\mathcal{F}}$ and $\nabla\mathcal{L}_{\mathcal{R}}$, respectively, the conflict can be quantified using the cosine similarity:
$cos(\nabla\mathcal{L}_{\mathcal{F}},\nabla\mathcal{L}_{\mathcal{R}})=\frac{\nabla\mathcal{L}_{\mathcal{F}}\cdot\nabla\mathcal{L}_{\mathcal{R}}}{\|\nabla\mathcal{L}_{\mathcal{F}}\|\cdot\|\nabla\mathcal{L}_{\mathcal{R}}\|}$ (12)
@@ 4.2.2 Orthogonalizing Gradient Conflict
$\nabla\mathcal{L}_{\mathcal{F}}^{\text{proj}}=\nabla\mathcal{L}_{\mathcal{F}}-\frac{\nabla\mathcal{L}_{\mathcal{F}}\cdot\nabla\mathcal{L}_{\mathcal{R}}}{\|\nabla\mathcal{L}_{\mathcal{R}}\|^{2}}\nabla\mathcal{L}_{\mathcal{R}}$ (13)
@@
$\nabla\mathcal{L}_{FALCON}=\alpha\nabla\mathcal{L}_{\mathcal{F}}^{\text{proj}}+\beta\nabla\mathcal{L}_{\mathcal{R}}$ (14)
@@
This mechanism mitigates gradient conflicts, enabling joint optimization while minimizing interference. By enforcing orthogonality between adjusted gradients, it approximates a Pareto-optimal solution. The model then updates its weights using the conflict-reduced gradient, allowing for more flexible adaptation. To further enhance efficiency and stability, we leverage the second-order optimizer Sophia (Liu et al., 2024a), as suggested in (Gu et al., 2024; Jia et al., 2024), for refined weight updates, ensuring a more effective and stable optimization process.
@@ 5 Experiments
To validate FALCON’s effectiveness, we conduct extensive experiments to answer the following research questions:
RQ1: Can MI guidance quantify the degree of knowledge entanglement, providing a measurable basis for parameter selection in unlearning?
RQ2: Does the FALCON outperform state-of-the-art train-time unlearning methods in terms of both unlearning effectiveness and utility?
RQ3: Can FALCON effectively resist recovery attempts of unlearned knowledge?
5.1 Experimental Setup
Datasets.
To answer the above research questions, we evaluate our method on four benchmark datasets that are highly relevant to unlearning effectiveness and utility performance: WMDP (Li et al., 2024), WikiText (Merity et al., 2016), and MMLU (Hendrycks et al., 2021a). The WMDP dataset evaluates the reduction of malicious use by targeting the removal of sensitive and harmful information. WikiText benchmarks model perplexity in general-purpose language understanding, while MMLU (Massive Multitask Language Understanding) is used to evaluate model utility across diverse domains after unlearning.
Models.
We test FALCON on three high-quality open-source pre-trained language models: Zephyr-7B-Beta (Tunstall et al., 2023), Yi-6B-Chat (Young et al., 2024), and Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), to evaluate the levels of forgetting and utility in LLMs after unlearning.
Baselines.
To ensure a fair comparison, we evaluate FALCON against several state-of-the-art training-time unlearning methods, including LLMU (Yao et al., 2023), SCRUB (Kurmanji et al., 2024), SSD (Foster et al., 2024), and RMU (Li et al., 2024). The details of each method are shown in Appendix D.
5.2 Mutual Information for Parameter Selection
@@
The heatmaps of MI shown in Figure 2 illustrate the degree of knowledge entanglement between the forget sets (WMDP-Bio, WMDP-Cyber) and the retained set (WikiText-2-raw-v1) across the layers of various LLMs. MI provides a quantitative and interpretable measure to identify layers where the representations of forget and retain datasets are least intertwined, offering a principled criterion for selecting suitable layers for unlearning.
Across all three models (Zephyr-7B-Beta, Mistral-7B-Instruct-v0.3, and Yi-6B-Chat), lower MI values are generally concentrated in the earlier layers, indicating more domain-specific and disentangled representations, which aligns with both intuition and experimental observations (Li et al., 2024). Notably, Yi-6B-Chat exhibits more complex entanglement patterns between biological and cybersecurity domains, presenting a greater difficulty for unlearning multi-domain knowledge and making it an ideal candidate for our effectiveness analysis experiments in Section 5.3. Moreover, MI-guided parameter selection can improve computational efficiency by narrowing the parameter optimization space compared to exhaustive methods like grid search. This approach scales effectively with increasing model complexity, providing a practical guiding indicator for unlearning in modern LLM architectures.
@@
We empirically evaluated MI’s effectiveness in guiding unlearning by analyzing gradient conflicts between forget and retain objectives across different layers in Mistral-7B-Instruct-v0.3. As shown in Figure 3, layers with low MI values demonstrate significantly reduced gradient conflicts, exhibiting cosine similarity values closer to zero. This indicates that these layers’ disentangled representations minimize interference between objectives, enabling more stable and efficient parameter updates. In contrast, layers with high MI values show pronounced conflicts through widely fluctuating and often negative cosine similarity values, reflecting the challenges posed by entangled representations.
Figure 3: Gradient conflicts across layers with minimum (blue) and maximum (orange) mutual information values computed during parameter selection.
These experimental results validate the use of MI as an effective guiding metric for unlearning optimization. Parameters identified with low MI values prove to be better targets for unlearning interventions, as they facilitate precise suppression of undesired knowledge while preserving critical information through reduced gradient conflicts. This empirical evidence strongly supports our strategy of using MI for parameter selection in the unlearning process.
5.3 Unlearning Effectiveness and Utility Analysis
Table 1: Performance comparison of unlearning effectiveness and utility across models and methods. Metrics with ($\uparrow$) indicate higher is better, while ($\downarrow$) indicate lower is better.
@@ 5.3 Unlearning Effectiveness and Utility Analysis
Table 2: Knowledge Recovery Results via Enhanced GCG Attack
@@
We evaluate FALCON against state-of-the-art unlearning methods across three LLM architectures shown in Table 1 and Appendix E.1, with our evaluation focusing on three key metrics: WMDP scores for measuring unlearning effectiveness, MMLU scores for assessing general knowledge retention, and perplexity (PPL) for model stability. Our primary objective is to minimize WMDP scores while maintaining MMLU and PPL values close to the base model’s performance (MMLU and PPL), as this indicates successful knowledge removal without compromising general capabilities. To ensure fair comparison, we prioritize maintaining general model utility and report each method’s best unlearning performance under this setting. The experimental results demonstrate FALCON’s superior performance compared to baseline methods, which often struggle to balance unlearning effectiveness with model utility and show increased uncertainty in their perplexity scores. On Zephyr-7B, FALCON achieves substantially lower scores on forgetting tasks while preserving the model’s general capabilities. This advantage becomes even more pronounced on Yi-6B-Chat, where knowledge representations are more deeply entangled: while RMU shows significant performance degradation in the biological domain when constrained to maintain MMLU score above 60%, FALCON maintains consistent unlearning effectiveness while achieving superior general task performance. Furthermore, FALCON demonstrates remarkable efficiency by achieving excellent results through single-layer intervention identified by mutual information, compared to RMU’s requirement of modifying at least three layers. These results validate the effectiveness of our fine-grained representation unalignment via contrastive and orthogonal projection mechanisms in achieving targeted unlearning while preserving model utility, even in challenging scenarios with complex knowledge entanglement.
5.4 Resistance Against Knowledge Recovery Attempts
Figure 4: Logit lens probing results on different components of Yi-6B-Chat.
We conduct extensive experiments on Yi-6B-Chat to evaluate FALCON’s resistance against knowledge recovery attempts (Łucki et al., 2024). Logit Lens (Patil et al., 2024), which projects intermediate activations onto the model’s vocabulary space, serves as a powerful technique for probing the model’s internal knowledge representations and potential recovery of unlearned information. As shown in Figure 4, the logit lens analysis across different architectural components such as MLP and attention layers demonstrates that the unlearned knowledge remains consistently inaccessible, with performance staying close to the unlearned baseline and far below the original model’s performance. Additionally, as shown in Table 2, FALCON exhibits strong resilience against Enhanced GCG in QA setting, a sophisticated prefix-optimization based jailbreaking attack (Thompson & Sklar, 2024). Even with increasing attack iterations, the recovered WMDP scores remain remarkably stable near the unlearned baseline, demonstrating that our approach achieves robust unlearning by fundamentally altering the model’s internal representations rather than merely achieve surface-level knowledge separation. Further evaluation using conversational templates for jailbreaking attacks (detailed in Appendix E.3) further validates our method’s robustness against knowledge recovery attempts. These results across both probing techniques validate FALCON’s effectiveness in creating a more permanent and recovery-resistant form of knowledge removal.
5.5 Ablation Study Analysis
To validate the effectiveness of FALCON’s components, we conduct ablation studies on Yi-6B-Chat. The baseline demonstrates a solid performance of 27.5% on WMDP and 60.3% on MMLU. Replacing the contrastive loss with MSE loss (w/o Loss) renders unlearning ineffective, emphasizing the necessity of the contrastive mechanism for precise knowledge separation. While removing gradient projection (w/o GP) or replacing POVs with random vectors (w/o POVs) has a minor impact on unlearning but degrades knowledge retention, highlighting their critical role in preserving model utility. These results empirically confirm that each component is essential for FALCON’s success in achieving precise unlearning while maintaining general model performance.
Table 3: Impact of component omission on performance.
Variant Omit WMDP ($\downarrow$) MMLU ($\uparrow$)
Baseline 27.5 60.3
w/o Loss 50.7 61.4
w/o GP 27.4 58.4
w/o POVs 27.6 57.6

### v2

@@ 4.2 Contrastive Orthogonal Unalignment
To achieve selective knowledge unlearning in LLMs, we first apply MI-guided parameter selection  [^2] Discussion on MI-guided parameter selection is shown in Appendix. F.2 to identify layers with minimal knowledge entanglement, which remains fixed throughout unlearning. We then devise Contrastive Orthogonal Unalignment through contrastive mechanisms and gradient projection, employing alternating strategy between forget and retain datasets to iteratively refine representations while balancing knowledge removal and retention objectives.
@@ 4.2.1 Contrastive Representation Unlearning
To facilitate thorough unlearning, we construct Principal Offset Vectors (POVs) that steer model activations away from undesired knowledge by redirecting updated forgettable representations into subspaces intentionally misaligned with the principal directions of frozen counterparts, as identified via SVD, thereby achieving representational decoupling within the model.
@@
$\small\mathcal{H}^{+}=\frac{f\left(r\cdot\left(I-w\sum_{i=1}^{K}v_{i}v_{i}^{\top}\right),\epsilon\right)}{\left|f\left(r\cdot\left(I-w\sum_{i=1}^{K}v_{i}v_{i}^{\top}\right),\epsilon\right)\right|}$ (8)
Here, $r\in\mathbb{R}^{D}$ is a randomly initialized vector, $w$ controls the influence of principal directions, and $I\in\mathbb{R}^{D\times D}$ is the identity matrix. The term $\epsilon$ introduces optional perturbations while $f(\cdot)$ is a flexible transformation operator, potentially including non-linear mappings (e.g., tanh), adaptive projections, or adversarially-inspired perturbations, enhancing disentanglement and recovery resistance. This design ensures $\mathcal{H}^{+}$ is directed away from dominant principal subspaces, combining deterministic guidance and transformations to improve robustness. Unlike generic random vectors, POVs deliberately target dominant features to improve adversarial robustness and unlearning efficacy.
For each input sample, we define three types of representations: the anchor representation $\mathcal{H}_{a}$ from the updated model for the forget set, the positive representation $\mathcal{H}^{+}$, given by the POV defined in Eq. (8), and the negative representations $\mathcal{H}^{-}$ from the frozen model. To ensure consistent scaling, all representations are normalized, and their similarity scores are measured using cosine similarity:
$S^{+}$ $=\sum_{d=1}^{D}\mathcal{H}_{a}[d]\cdot\mathcal{H}^{+}[d],\qquad S^{-}=\sum_{z=1}^{\mathcal{Z}}\sum_{d=1}^{D}\mathcal{H}_{a}[d]\cdot\mathcal{H}_{z}^{-}[d]$ (9)
@@
$\small\mathcal{L}_{\mathcal{F}}=-\frac{1}{|B|}\sum_{b=1}^{|B|}\log\frac{\exp(S_{b}^{+}/\tau)}{\exp(S_{b}^{+}/\tau)+\sum_{b=1}^{\mathcal{N}}\exp(S_{b}^{-}/\tau)}$ (10)
where $\tau$ is a temperature scaling parameter. This loss encourages the updated model’s representations to align with the POVs while diverging from the frozen model’s representations of undesired knowledge. By leveraging both directional guidance through POVs and contrastive learning, our approach achieves more precise and efficient representation unalignment in activation space.
In addition to unlearning undesired representations, preserving critical knowledge for downstream tasks is essential. We define a retain loss in Eq. (11) to measure alignment between the updated model’s activations ($\mathcal{H}^{u}$) and frozen model’s activations ($\mathcal{H}^{f}$) for the retain set. This retention alignment loss, functioning as a self-supervised variant of contrastive loss, maximizes consistency between updated and frozen activations to ensure effective knowledge preservation during unlearning.
@@
4.2.2 Orthogonalizing Gradient for Conflict Resolution
After computing the forget loss $\mathcal{L}_{\mathcal{F}}$ and retain loss $\mathcal{L}_{\mathcal{R}}$, we address optimization direction misalignment between unlearning and retaining by employing a gradient projection mechanism that orthogonalizes conflicting gradients onto subspaces, minimizing interference and promoting balanced optimization. Given the gradients of the forget and retain losses, denoted as $\nabla\mathcal{L}_{\mathcal{F}}$ and $\nabla\mathcal{L}_{\mathcal{R}}$, respectively, the conflict can be quantified using the cosine similarity:
$\small cos(\nabla\mathcal{L}_{\mathcal{F}},\nabla\mathcal{L}_{\mathcal{R}})=\frac{\nabla\mathcal{L}_{\mathcal{F}}\cdot\nabla\mathcal{L}_{\mathcal{R}}}$$\|\nabla\mathcal{L}_{\mathcal{F}}\|\cdot\|\nabla\mathcal{L}$$_{\mathcal{R}}\|$} (12)
@@ 4.2.2 Orthogonalizing Gradient for Conflict Resolution
$\small\nabla\mathcal{L}_{\mathcal{F}}^{\text{proj}}=\nabla\mathcal{L}_{\mathcal{F}}-\frac{\nabla\mathcal{L}_{\mathcal{F}}\cdot\nabla\mathcal{L}_{\mathcal{R}}}{\|\nabla\mathcal{L}_{\mathcal{R}}\|^{2}}\nabla\mathcal{L}_{\mathcal{R}}$ (13)
@@
$\small\nabla\mathcal{L}_{FALCON}=\alpha\nabla\mathcal{L}_{\mathcal{F}}^{\text{proj}}+\beta\nabla\mathcal{L}_{\mathcal{R}}$ (14)
@@
This mechanism mitigates gradient conflicts, enabling joint optimization while minimizing interference. By enforcing orthogonality between adjusted gradients, it approximates a Pareto-optimal solution. The model then updates its weights using the conflict-reduced gradient, allowing for more flexible adaptation. To further enhance efficiency and stability, we leverage the second-order optimizer Sophia liu2024sophia , as suggested in gu2024second ; jia2024soul , for refined weight updates, ensuring a more effective and stable optimization process for selective knowledge unlearning.
@@ 5 Experiments
To validate FALCON’s effectiveness, we conduct extensive experiments to answer the following research questions: RQ1: Does FALCON with MI guidance, establish a quantifiable measure for principled parameter selection while achieving superior performance in harmful knowledge unlearning tasks? (Section 5.1) RQ2: Does FALCON maintain strong generalizability across diverse unlearning tasks including entity unlearning and copyrighted content unlearning? (Section 5.2) RQ3: Beyond efficient parameter space reduction through MI guidance, does FALCON’s algorithmic design offer competitive computational efficiency? (Appendix E.3) RQ4: Can FALCON effectively resist recovery attempts of unlearned knowledge? (Section 5.3). More complete experiments and ablation study are shown in Appendix E.
5.1 Harmful Knowledge Unlearning
To validate RQ1, we use the WMDP li2024the benchmark for harmful knowledge unlearning assessment, WikiText merity2016pointer for measuring perplexity, and MMLU hendrycks2021measuring for evaluating model utility. We test FALCON on three pre-trained LLMs: Zephyr-7B-Beta tunstall2023zephyr , Yi-6B-Chat young2024yi , and Mistral-7B-Instruct-v0.3 jiang2023mistral , comparing against all baselines from li2024the , with details in Appendix D.
5.1.1 Mutual Information for Parameter Selection
@@
Figure 2 presents MI heatmaps illustrating knowledge entanglement between forget sets (WMDP-Bio, WMDP-Cyber) and the retain set (WikiText-2-raw-v1) across LLM layers. This metric provides an interpretable measure for identifying layers with minimal entanglement for targeted unlearning. All models show lower MI values in earlier layers, indicating more domain-specific and disentangled representations, which aligns with both intuition and experimental observations li2024the . Yi-6B-Chat demonstrates particularly complex entanglement patterns between domains, presenting a greater difficulty for unlearning multi-domain knowledge and making it an ideal candidate for our effectiveness analysis experiments in Section 5.1.2. Beyond identifying optimal intervention parameters, MI-guided selection improves efficiency by narrowing the parameter search space compared to exhaustive methods like grid search, scaling effectively with model complexity.
@@
We empirically validate the underlying principle of MI
Figure 3: Gradient conflicts across layers with minimum (blue) and maximum (orange) MI values computed during parameter selection in Mistral-7B.
guidance by analyzing gradient conflicts between forget and retain objectives across layers. As shown in Figure 3, layers with low MI values exhibit significantly reduced conflicts, with cosine similarities near zero, indicating minimal interference between objectives. Conversely, high-MI layers show pronounced, fluctuating conflicts, highlighting the issues of entangled representations. These results confirm that mutual information is a reliable auxiliary signal for guiding parameter selection, as low-MI parameters reduce interference, support stable updates, and help mitigate conflicts between unlearning and retention goals.
5.1.2 Unlearning Effectiveness and Utility Analysis
Table 1: Unlearning effectiveness and utility across models and methods. Metrics with ($\uparrow$) indicate preferable increases; ($\downarrow$) indicate preferable decreases.
@@ 5.1.2 Unlearning Effectiveness and Utility Analysis
We evaluate FALCON against all baseline methods across three LLM architectures shown in Table 1 and Appendix E.1, with our evaluation focusing on three key metrics: WMDP scores for measuring unlearning effectiveness, MMLU scores for assessing general knowledge retention, and perplexity (PPL) for model stability. Our primary objective is to minimize WMDP scores while maintaining MMLU and PPL values close to the base model’s performance (MMLU and PPL), as this indicates successful knowledge removal without compromising general capabilities. To ensure quantifiable comparison, we prioritize maintaining general model utility and report each method’s best unlearning performance under this setting. Results demonstrate FALCON’s superior performance compared to baselines that struggle with effectiveness-utility balance and show increased uncertainty in their perplexity. On Zephyr-7B, FALCON achieve lower forgetting scores while preserving general capabilities. This advantage is more clear on Yi-6B-Chat with its complex knowledge entanglement: RMU show significant biological domain degradation when constrained to maintain MMLU above 60%, while FALCON maintain consistent effectiveness with superior general performance. These findings validate our fine-grained representation-guided mechanisms for targeted unlearning with preserved utility, even in scenarios with complex knowledge entanglement.
5.2 Cross-Domain Generalizability Assessment
To address RQ2, we conduct additional experiments on copyrighted content and entity unlearning using the MUSE shi2024muse and TOFU maini2024tofu benchmarks with additional baselines openunlearning2025 . For RQ3, we compare computational efficiency across methods in Appendix E.3). All aforementioned experiments utilize first-order optimizers for fair comparison, with complete implementation details in Appendix D.
5.2.1 Copyrighted Content Unlearning
For copyrighted content unlearning, we utilize the MUSE benchmark and Llama-2-7b-hf to assess FALCON’s effectiveness in removing protected news articles while preserving general capabilities. As shown in Table 2, FALCON achieved the lowest forget metrics scores (0.02 and 0.03) while maintaining competitive retention (0.54). Unlike baselines, FALCON consistently balanced copyright removal with knowledge preservation, demonstrating broader applicability beyond harmful content removal.
Table 2: Evaluation on MUSE News over 10 epochs.
[…]
Finetuned 0.64 0.58 0.55
Retain 0.33 0.21 0.56
GradAscent 0.00 0.00 0.00
GradDiff 0.41 8.92e-3 0.37
NPO 0.56 0.35 0.51
SimNPO 0.54 0.36 0.51
RMU 0.48 0.05 0.51
FALCON 0.02 0.03 0.54
5.2.2 Entity Unlearning
Table 3: TOFU evaluation across varying sizes over 10 epochs.
Method Forget01 Forget05 Forget10
FQ MU FQ MU FQ MU
Finetuned 0.01 0.60 2.96e-13 0.60 8.08e-22 0.6
Retain 1.0 0.60 1.0 0.60 1.0 0.59
GradAscent 0.27 0.33 1.94e-119 0 1.06e-239 0
GradDiff 0.77 0.43 2.04e-110 0.22 1.06e-239 0.49
IdkDPO 0.01 0.51 4.02e-06 0.04 4.26e-10 0.08
NPO 0.92 0.56 0.32 0.42 0.02 0.46
RMU 0.16 0.55 1.46e-7 0.57 1.4e-20 0.59
FALCON 0.99 0.55 0.92 0.59 0.52 0.60
[…]
5.3 Resistance Against Knowledge Recovery Attempts
Figure 4: Logit lens probing results on different components.
We conduct experiments on Yi-6B-Chat to evaluate FALCON’s resistance against knowledge recovery attempts ucki2024an for RQ4. Logit Lens patil2024can , which projects intermediate activations onto the model’s vocabulary space, serves as a powerful technique for probing the model’s internal knowledge representations and potential recovery of unlearned information. As shown in Figure 4, the logit lens analysis across different architectural components such as MLP and attention layers demonstrates that the unlearned knowledge remains consistently inaccessible, with performance staying close to the unlearned baseline and far below the original model’s performance.
Additionally, as shown in Table 4, FALCON exhibits strong resilience against enhanced GCG in QA setting, an advanced prefix-optimization based jailbreaking attack that compromises other baselines such as RMU thompson2024flrt . Even with increasing attack iterations, the recovered WMDP scores remain close to the unlearned baseline, demonstrating robust unlearning through fundamental changes to the model’s internal representations rather than superficial knowledge masking. Further evaluation using conversational templates for jailbreaking attacks (detailed in Appendix E.6) further validates our method’s robustness against knowledge recovery attempts.
Table 4: Knowledge recovery results via enhanced GCG attack.
@@ 5.3 Resistance Against Knowledge Recovery Attempts
These results across both probing techniques validate FALCON’s effectiveness in creating a more permanent and recovery-resistant form of knowledge removal.
6 Practical Implications of LLM Unlearning for Responsible AI
The problem setting addressed by FALCON [^3] Additional discussions are provided in Appendix F. stems from the growing challenge of directly employing LLMs or deploying them as autonomous agents in safety-critical environments HU2025103779 ; hu2025tapasfreetrainingfreeadaptation . As these models become increasingly embedded in diverse real-world applications, selectively removing undesired or harmful knowledge after deployment remains difficult liu2024rethinking . Unlike conventional machine learning models where unwanted data can simply be excluded in future training cycles, LLMs encode information across billions of parameters, making precise removal extremely challenging. This limitation creates a critical gap between learning capabilities and responsible deployment. The issue is further amplified by regulatory demands such as the GDPR’s “right to be forgotten” regulation2016regulation , and by empirical evidence that even state-of-the-art LLMs and their agentic variants can inadvertently reproduce sensitive or hazardous content when prompted, raising urgent concerns about information safety and controllability.
FALCON provides a fine-grained unlearning mechanism that identifies harmful knowledge and decouples it from beneficial reasoning. This targeted process enables models to forget unsafe information while retaining legitimate competence, supporting the emerging need for responsible LLM deployment ibm2024forget ; hu2025positionresponsiblellmempoweredmultiagent . As LLMs operate in dynamic, real-world contexts, the capacity for precise and interpretable knowledge modification becomes essential for responsible AI. We advocate viewing unlearning not as an academic objective but as core practical infrastructure for transparent, compliant, and responsible AI systems.